show that cosx(270degree +theta)/cos(360degree - theta)= tantheta
1
Expert's answer
2020-10-11T18:38:59-0400

cos(270o+θ)/cos(360o-θ)=(cos(270o)cos(θ)-sin(270o)sin(θ))/(cos(360o)cos(θ)+sin(360o)sin(θ))=(0*cos(θ)-(-1)*sin(θ))/(1*cos(θ)+0*sin(θ))=sin(θ)/cos(θ)=tan(θ)

Q.E.D.